Please answer the following

QUESTION 1

  1. If a 1.5% cash discount is allowed for early payment, the amount of the discount on a $2,000 invoice would be:

    A.

    $1,970

    B.

    $300

    C.

    $30

    D.

    $1.50

20 points   

QUESTION 2

  1. If A = L + S, what is S if A is $54,000 and L is $23,000?

    A.

    $21,000

    B.

    $31,000

    C.

    $77,000

    D.

    $27,000

20 points   

QUESTION 3

  1. Ed gets paid $4,500 per month salary plus 4% commission on all sales greater than $50,000.  If he sold $72,000 last month, what was his total gross pay?

    A.

    $4,500

    B.

    $5,380

    C.

    $7,380

    D.

    $2,880

20 points   

QUESTION 4

  1. Real estate taxes are based on the assessed value of the property. Calculate the tax due on a house with an assessed value of $160,000. The tax rate is 6.255% of the assessed value.

    A.

    $10,008

    B.

    $9,600

    C.

    $18,000

    D.

    $10,800

20 points   

QUESTION 5

  1. The simple interest formula for calculating interest for less than one year is: Interest = Principal x Rate x Term/360. Solve for the Principal if the Interest is $120; Rate is 3%; Term is 60 days.

    A.

    $2,160

    B.

    $21,600

    C.

    $12,000

    D.

    $24,000
Answer #1

1. Option C

Discount on invoice amount = 2000*1.5% = 30

2. Option B

S = A- L = 54,000-23,000 = 31,000

3. Option B

Gross pay = 4,500 + (72000-50000)*4% = 5380

4. Option A

Tax due = 160,000*6.255% = 10,008

5. Option D

Principal = I/RT = 120/3%*60/360 =24,000